Louis sold his farm during the current taxable year. At the date of the sale, the farm had an adjusted basis of $212,000 and was encumbered by a mortgage of $190,000. The buyer paid him $110,000 in cash, agreed to take the title subject to the $190,000 mortgage, and agreed to pay him $80,000 with interest at 9% one year from the date of sale. How much is Louis' recognized gain on the sale?
1
Expert's answer
2014-06-18T09:30:25-0400
Total selling price means the amount of cash, the assumed note, and the installment note.
Thus, selling price = 190,000+110,000+80,000 = 380,000.
Realized gain is the amount of money actually earned in the sale of an asset.
RealG = 380,000 - 212,000 = 168,000.
Recognized gain is a profit earned from the sale of an asset.
RecG = 168,000 * 300,000 / 380,000 = 132,632. The balance of the gain will be recognized when the amount of 80,000 is collected.